Suppose you toss two 6 sided dice on a table. What is the most likely number...

  1. Suppose you toss two 6 sided dice on a table.
    1. What is the most likely number to roll?
    2. Why?
    3. What is the probability of rolling a total of 4?
Answer #1

a)most likely number to roll =7

b)as a sum of 7 appears in maximum number of ways =6 {(1,6).(2.5),(3,4),(4,3),(5,2),(6,1)}

c)

total number of outcomes =6*6=36

number of ways getting a 4 as sum =3 {(1,3),(2,2),(3,1)}

hence probability of rolling a total of 4 =3/36 =1/12